OpenStudy (anonymous):

i got this wrong in my homework i was wondering how thy get this answear Determine the percent from two numbers given here. Fifty is what percent of 195? which was 30% does that mean i have to do these steps - finding the difference between the two numbers (subtraction) - writing the answer as a fraction over the larger no. - multiplying it by 100 (percentage)

OpenStudy (debbieg):

I don't really understand your question... if the question is, "Fifty is what percent of 195".... you can always set it up: \[\Large \frac{ 50 }{ 195 }=\frac{ p }{ 100 }\] so: \[\Large \frac{ 50 }{ 195 }\cdot100=p\] That is, just do the division 50/195, and multiply that by 100 to convert the decimal to a percent. But, 50 is NOT 30% of 195, so I'm not sure what that part of your question means??

OpenStudy (anonymous):

yea thats my question but my homework is online and when i got it wrong it said the answer was 30percent and i been trying to figure out how they got 30 which ur probably right

OpenStudy (debbieg):

(50/195) * 100 = approx 25.6% Probably a typo in the problem/answer.... lol.

OpenStudy (anonymous):

Lol I hope thanks tho